A genius that directs all their evil towards art creates a trap: A painting owner weighing 90 kg falls from a height of 50 m onto a soft mat, which takes all of their kinetic energy into a spinning circular saw that is set up just touching a painting, with a mass of 10 kg and a radius of ½ m. How fast does the saw spin after the person’s fall?
Answer #1

here,

height ,h = 50 m

mass , m = 90 kg

let the final speed of saw be v

using conservation of energy

0.5 * I * w^2 = M * g * h

0.5 * ( 0.5 * m * r^2) * ( v/r)^2 = M * g * h

0.5 * ( 0.5 * 10) * ( v^2) = 90 * 9.81 * 50

solving for v

v = 132.9 m/s

the speed of saw is 132.9 m/s